OpenStudy (anonymous):

if a roll of a certain kind of wrapping paper is unrolled, the wrapping paper forms a rectangle that is 3 ft wide and 20 ft long. The wrapping paper is to be covered with an array of ladyybugs by repeating the design show. (the design is 3 inches long by 2 inches wide, there are 12 ladybugs in this dimension) How many ladybugs will be on the wrapping paper?

OpenStudy (anonymous):

3/12 * 2/12 = 6/144 ft^2 3*20 = 60 ft ^2 60/ (6/144) =1440